A recent survey found that 68.6% of people own their homes. In a random sample of 150 heads of household, 92 responded that they own their homes. At a=0.01, does that suggest a difference for the proportion? Find the CV and compute the test value (TV) , and state your conclusion.

 


1
Expert's answer
2021-05-07T09:28:28-0400

H0:p=0.686

H1:p≠0.686 (claim)


"p=0.686, n=150, \hat{p}=92 /150=0.613"


"CV \text{ at } \alpha=0.01 \text{ is equal to }2.58"


"z=\dfrac{(\hat{p}-p)}{\sqrt{\frac{p(1-p)}{N}}}" "=\\dfrac{(0.613-0.686)}{\\sqrt{0.001436}} \n\u200b\t\n =-1.918"


p-value"= (2)P(z < -1.918) = (2)(0.0276) = 0.0552 (\text{ two-tailed test })"


Since p-value > 0.01, do NOT reject the null hypothesis. There is significant evidence that the mean proportion of homeowners is not different from 68.6%.